OpenStudy (anonymous):

Find the simple interest if the principal is $4,000, the rate is 5 1/2%, and the investment period is 30 months.

OpenStudy (anonymous):

A=P(1+r/n)^(nt) A = final amount P = principal/initial amount r = annual rate of interest n = number of times the interest is compounded per year t = number of years A=4000(1+0.055/12)^(12*2.5) A=4588.17 4588.17-4000=$588.17 interest

OpenStudy (anonymous):

The answer is $550 but I'm still not sure how they got this. :)

OpenStudy (rogue):

EnchantixPixie's n is wrong. It is supposed to be n = 12/30 = 0.4 times a year, not n = 30/12 = 2.5 times a year.